An echo is heard in 3s.what is the distance of the reflecting surface from the source,given that the speed of sound is 342m/s
1
Expert's answer
2020-05-26T12:52:34-0400
"d=\frac{vt}{2}"

"d=\frac{342\cdot 3}{2}=513 \ m"